I am greater than 50. The sum of my digits is 10 and the difference between my figures is 2. What number am I if I am a two digit number
Answer by macston(5194)

Let one of the digits be X, let the other be Y
X+Y=10
X-Y=2; X=Y+2 Substitute this for X
(Y+2)+Y=10
2Y=8
Y=4 One of the digits is 4
X+4=10 Subtract 4 from each side
X=6 One of the digits is six
The combination of these two digits that form a number greater than 50 is 64:
ANSWER: The number is 64.